### v2

# Surf-NeRF: Surface Regularised Neural Radiance Fields

Abstract
Neural Radiance Fields (NeRFs) provide a high fidelity, continuous scene representation that can realistically represent complex behaviour of light. Despite works like Ref-NeRF improving geometry through physics-inspired models, the ability for a NeRF to overcome shape-radiance ambiguity and converge to a representation consistent with real geometry remains limited. We demonstrate how both curriculum learning of a surface light field model and using a lattice-based hash encoding helps a NeRF converge towards a more geometrically accurate scene representation. We introduce four regularisation terms to impose geometric smoothness, consistency of normals, and a separation of Lambertian and specular appearance at geometry in the scene, conforming to physical models. Our approach yields 28% more accurate normals than traditional grid-based NeRF variants with reflection parameterisation. Our approach more accurately separates view-dependent appearance, conditioning a NeRF to have a geometric representation consistent with the captured scene. We demonstrate compatibility of our method with existing NeRF variants, as a key step in enabling radiance-based representations for geometry critical applications.
1 Introduction
Neural Radiance Fields (NeRFs) [34] provide an efficient coordinate based scene representation with wide applications to computer vision, robotics and beyond. The ray-based volumetric rendering formulation produces realistic novel views of a scene, encompassing view-dependent scene appearance. Complex appearances like specularity, transparency and interreflection lead to shape-radiance ambiguity where scene geometry is not uniquely represented, often producing imagined geometries.
Figure 1: Surf-NeRF uses a permutohedral lattice encoding and the properties of surface light fields to regularise local regions of a NeRF. This improves consistency and smoothness in these regions by accumulating density at a surface, leading to improved geometry and more physically viable appearance components (ours, right) compared to current state-of-the-art ([43], left). Changing the frequency of regularisation balances between geometric fidelity and appearance separation.
Re-parameterisation of the radiance in the scene as in Ref-NeRF [43] has shown improvements over previous state-of-the-art variants, however there exists an incomplete separation of the scene’s Lambertian appearance from the specular and no explicit constraint on the placement of density. It is still common for the NeRF to place regions of density behind surfaces or in front of the camera as detached geometry (floaters) to explain complex phenomena. This leads to a non-realistic geometric scene representation, as in Figure 1.
Applications like robotics and 3D modelling require both accurate geometry and realistic appearance. Poor representation of physical geometry currently poses a large hurdle to the widespread adoption of radiance-based scene representation for geometry critical tasks. Robotic manipulation, autonomous navigation and mapping require an accurate scene structure to minimise the gap between a representation and the real world, for example when grasping metal objects or navigating near reflective windows. Similarly, traditional structure-from-motion pipelines have great difficulty in reconstructing visually complex objects including reflective and transparent surfaces.
We address geometric inaccuracy arising from view-dependent phenomena with the insight that there are multiple formulations of the plenoptic function which can produce viable scene representations. A surface light field [48] describes the plenoptic function as light originating from a geometrically smooth surface, which we encode using a permutohedral lattice hash encoding. Our insight lies in that we can push the NeRF towards a more geometrically accurate representation in the same rendering framework using additional geometric and appearance based regularisation via curriculum learning. We use a first surface assumption to describe the location of geometry in the scene.
Using a second sampling of the NeRF at these points, we regularise density to produce smoothly-varying normals and thin, continuous sheets of density which more realistically represent scene geometry. Regularising the view-dependent appearance to remove Lambertian components we reduce shape-radiance ambiguity and encouraging more correct structure of the surface. In summary, we make the following contributions:
  1. 1.
    We devise a novel regularisation approach which uses the structure of a neural radiance field to sample density, normals and appearance in the vicinity of geometry in the scene, allowing for additional representation-driven regularisation terms to be applied.
  2. 2.
    We apply local regularisation consistent with a surface light field radiance model, including geometric smoothness of density, local consistency of normals and a physically correct separation of Lambertian and specular appearance using a light interaction model.
  3. 3.
    We introduce a permutohedral lattice encoding for NeRF which better represents non-planar and complex geometry over traditional cubic hash grids.
  4. 4.
    We introduce the Koala dataset of four scenes with visually challenging objects and complex geometry to benchmark our approach.
We benchmark our approach on state-of-the-art physics based NeRF variants, however our methodology may also be applied to other NeRF frameworks and likely adapted to other radiance based representations.
This work is a key step in the deployment of NeRFs as a scene representation where both geometric and visual fidelity are critical, like robotic manipulation and navigation in complex unstructured environments.
2 Related Work
Neural Scene Representations: Implicit neural field approaches [25] produce continuous and high visual fidelity scene representations. These can be queried anywhere within the training set, balancing visual and geometric fidelity. In this work, we improve the geometry of a NeRF and maintain visual fidelity, reducing reliance on introduced geometry to meet scene appearance.
NeRFs [34] leverage the efficiency of ray-based construction similar to light fields with a volumetric scene representation of points which emit light. Subsequent works improved the fidelity of representation [4, 5, 6], recovery of a static scene [40, 42, 33] and performance around view-dependent phenomena such as reflections [43, 23]. However, relying on a purely volumetric rendering approach allows the network to imagine geometries to explain view-dependent phenomena, particularly in cases where regions of the scene are underconstrained [38, 28, 36] or where light does not follow the physical model of a straight ray through the scene [8]. Additional regularisation terms can dramatically improve the quality of rendering in these cases [36, 28], e.g. improving the accuracy of normals significantly helps representation of complex appearances [32, 43].
More recent works [44] have shown that ray-tracing multiple reflection rays can improve representation reflective scene element, albeit with extremely high computation cost. In this work we use properties of surface light fields to reduce the reliance of NeRFs on additional geometries, maintaining a single-pass volumetric rendering approach.
Radiance Based Scenes: Recent advances in explicit representations like 3D Gaussian splatting [27] (3DGS) have demonstrated improved rendering quality over volumetric implicit models, reflective objects rely heavily on introduced Gaussians behind surfaces to represent view dependence. Although subsequent work has sought to improve the geometric fidelity of the resulting representation [24, 26, 29, 47, 46], these involve substantial changes to the underlying representation or simplified geometry models due to the sparsity of the explicit radiance field. Our approach derives key benefits from the continuity of implicit radiance fields to formulate regularisation terms. We anticipate similarly inspired terms might be formulated for explicit methods like 3DGS, Radiant Foam and others [27, 22, 31, 24].
Surface Representations: Signed distance fields (SDFs) have garnered significant attention [51, 52, 16, 18, 3, 37] as they can represent smooth and continuous geometries in space and may be generated from multi-view constraints alone. For diffuse planar objects, this representation provides a smooth and high-quality reconstruction; however, more highly view-dependent appearance and complex geometry (thin structures, concave geometries) are not well represented [45, 39, 21]. Alternate volumetric encodings better suited to complex geometry like PermutoSDF [39] have substantially improved this representation, which we leverage in this work. Some works have included a reflection parameterisation [21], however, struggle to reconstruct 3D scenes with fine detail, thin structures, and sharp changes compared to volumetric scenes. Appearance and geometry have an intrinsic relationship within neural representations, with geometry being altered to account for non-Lambertian appearance [18]. We maintain a volumetric representation allowing for complex geometry, but with separated view-dependent and -independent appearance.
Appearance Vs. Geometry: Inverse rendering with mesh-based models produces a definite [20] separation of the scene’s appearance, geometry and environment, which are necessary to create realistic renderings of scenes under new conditions. Performing this without prior knowledge of the scene proves to be an immensely difficult task [53, 12, 14, 50], however, utilising physically-based rendering fused with learnt appearance [10, 11] provides sufficient constraints to recover the components of the scene.
Radiance field approaches seeking to learn an accurate scene representation under few view scenarios have used depth consistency [41], additional depth supervision [56] and regularisation using priors [36]. Incorporating an understanding of the physical interactions of light within a scene and its effect on appearance [54] under a solid surface assumption enables accurate geometry to be learnt alongside appearance. In the presence of specularities and other visual phenomena it is difficult to disentangle where appearance has been baked into geometry [49].
Without the need for recovering a bidirectional reflectance distribution function (BRDF), light field approaches [19] provide a more generalised framework. This enables a clear separation between appearance and geometry, reducing the need to acquire environmental view-dependent effects. Our work leverages similar light-field characteristics to recover geometry more accurately within the volumetric framework of NeRF, adding a prior to how geometry and appearance should present in the scene.
3 Method
Figure 2: An overview of the Surf-NeRF methodology. We use a first surface assumption of light to locate samples $\mathbf{x}_{\star}$ which are likely to lie on a surface in the scene during regular training steps. Sampling at the surface in multiple directions, and at nearby points drawn from a sphere $\mathbf{p}$, we impose regularisation on directional and spatial behaviour in this region. We separate the Lambertian component from the specular colour channel $\mathbf{c}_{s}$ by using a total variation prior. Using neighbouring points we also regularise geometric smoothness on density $\tau$ and consistency of normals $\hat{\mathbf{n}}$ leading to improved, continuous geometry.
Surf-NeRF introduces novel regularisation to locally enforce the properties of surface light fields. By regularising towards this representation, we represent smooth geometry with a physically viable separation of appearance and geometry producing more geometrically accurate radiance fields. This reduces shape radiance ambiguity by encouraging continuous regions of density with a smoothly varying view-dependent appearance. A visual depiction of our methodology is shown in Figure 2.
3.1 Preliminaries
A surface light field is a subset of the plenoptic function [9] that provides the colour of a light ray originating from a surface within the scene. Solid scene geometries are well represented given the decoupled parameterisation of the scene geometry and radiance on these surfaces.
A surface light field $L$ exists strictly on a surface geometry $G$ mapping directions in the unit 2-sphere ($\mathbf{S}^{2}$) to radiance, $L\,:\,G\times\mathbf{S}^{2}\rightarrow\mathbf{c}$, for an RGB colour triplet $\mathbf{c}$ [48].
Few real surfaces have an entirely view-dependent appearance. Similar to traditional light fields [19], a surface light field may be decomposed into a Lambertian (diffuse) reflectance and a view-dependent component [48, 30]. A surface light field has a strong assumption that the radiance seen from a given direction is provided entirely from a piecewise smooth geometry in the scene. This accordingly models phenomena such as volumetric scattering, reflection or transmission as a function of surface position rather than in a volumetric (NeRF) or physics-based (rendering engine) manner. In this way, a surface light field is decoupled from the geometry of a scene, meaning a surface may be deformed while its appearance is maintained [48]. Our approach is motivated by this decoupling to encourage Lambertian radiance to exist on smooth sheets of density, or surfaces, with a view-dependent colour.
3.2 Model
Reflection Parameterisation: Our method builds on the reflection parameterisation in Ref-NeRF [43], which splits the scene into a diffuse and specular appearance term similar to a surface light field. This parameterisation provides enhanced results across Lambertian and specular scenes over original NeRF variants [34, 4] by learning a spatially varying diffuse colour $\mathbf{c}_{d}$, specular tint $\mathbf{s}$, rendering normal $\hat{\mathbf{n}}^{\prime}$ and a view-dependent specular colour $\mathbf{c}_{s}$ for each point in the scene. The final colour of a ray in this parameterisation is given as $\mathbf{c}=\mathbf{c}_{d}+\mathbf{s}\mathbf{c}_{s}$, a linear combination of the diffuse and specular terms. Ref-NeRF struggles to entirely separate Lambertian and specular appearance, utilising density instead to explain complex phenomena such as non-planar reflection, anisotropic reflection and interreflection, leading to the results seen in Figure 1.
Permutohedral Lattice Hash Encoding: We replace the regular cubic hash grid used in prior NeRF variants [7, 35] with a hash-based permutohedral lattice [1]. Prior work [39], has used this structure for its memory efficiency, but our insight lies in the volumetric coverage and relative density of simplices in this structure. A permutohedral lattice, is the most efficient lattice solution the sphere covering problem [15] in low dimensions. Vertices in this lattice are arranged evenly throughout $d$-space and packed more densely [1], allowing for more accurate interpolation of features over cubic volume elements. In $d=3$-space, simplices take the form of regular tetrahedra which can more accurately represent complex non-planar geometries and closely approximating the true geometry of the scene. Our permutohedral lattice encoding therefore has two main benefits over cubic grids: 1) half the memory accesses per interpolation by using barycentric interpolation of a tetrahedron [39], and 2) dramatically improved representation of non-planar surfaces derived from density by representing curved geometry with multi-resolution tetrahedral features.
We apply our regularisation losses denoted by $\mathcal{L}$ and permutohedral lattice encoding to ZipNeRF [6] using the Ref-NeRF parameterisation, leveraging its state of the art performance with grid-based encoding [35]. We make no modifications to the model itself beyond our hash encoding, but include a second, data-driven sampling to impose physically-inspired regularisation. We maintain two proposal networks and one NeRF network with 64 and 32 samples per ray. Importantly, our regularisation terms are extensible to other NeRF variants, and may be applied after the main training as a fine-tuning stage. An overview of model structure is provided in the supplementary materials.
3.3 Sampling Radiance at a Surface
We regularise the surface and its light field by sampling a batch of positions and directions at the point where the ray intersects a surface in the scene, as shown in Figure 3. These local regularisation terms are based on surface light field properties, current scene geometry and appearance at this location.
Where prior works have sampled unseen image patches [36] to encourage consistent depths, we sample a batch of unseen rays localised at a surface point in the scene to encourage local geometric continuity. This batch approach also has significant benefit over single perturbed points seen in prior work [37, 55], as it allows for changes in surface orientation and structure not captured by a single sample to be accounted for, as we detail in Section 3.4.
We utilise a first-surface assumption to infer the location of geometry; we assume the majority of radiance is emitted by the point of highest density closest to the camera. Our candidate surface is the first point along a ray with weight greater than the median weight of the ray, preventing regularisation from occurring behind the true location of the surface. This minimises sampling around points which are occluded and therefore not well positioned with multi-view constraints. Choosing the median ensures that this selection is more robust to skewed weight distributions, particularly early during training. This is shown in Figure 2 left.
Using the point $\mathbf{x}_{\star}$, ray origin $\mathbf{o}$ and direction $\mathbf{d}$, surface normal $\mathbf{\hat{n}}_{\star}$, and covariance $\mathbf{\Sigma}_{\star}$ of this sample’s integrated positional encoding [4], we generate two new batches: a spatial batch to regularise density and a directional batch to regularise appearance. Spatial samples query the density field $\tau$ in the local 3D region of the surface at $\mathbf{x}_{\star}$ and the directional samples query the distribution of view-dependent colour $\mathbf{c}_{s}$ at different viewing angles.
We adapt a deterministic sampling scheme on the sphere [17] to produce uniformly distributed points used in both regularisation batches. Samples drawn from a von Mises-Fischer distribution with concentration parameter $\kappa=0$ are uniformly distributed on the unit 2-sphere $\mathbf{S}^{2}\subset\mathbb{R}^{3}$. We sample this distribution deterministically using a Fibonacci-Kronecker lattice as proposed by [17] for $N$ samples. This is partitioned into $\log_{2}N$ shells sampling the unit ball. Similar to the unscented sampling presented in ZipNeRF [6], we apply a random rotation [2] to these samples during training to avoid any bias from the orientation of the sampling scheme, arriving at samples $\mathbf{p}$. Further details are provided in the supplementary material.
We construct both batches of samples from virtual rays passing through this sphere centred at the surface. The points $p$ define directions $\mathbf{d}_{s}=\hat{\mathbf{p}}$, origins $\mathbf{o}_{s}$ and covariances $\mathbf{\Sigma}_{s}$ to construct new conical frusta. The origins and covariances for these rays,
$\mathbf{o}_{s}=\mathbf{R}_{s}(\mathbf{o}-\mathbf{x}_{\star})+\mathbf{x}_{\star},\quad\mathbf{\Sigma}_{s}=\mathbf{R}_{s}\mathbf{\Sigma}_{\star}\mathbf{R}_{s}^{\top},$ (1a,b)
are defined by the rotation matrix $\mathbf{R}_{s}$ which rotates the original ray direction $\mathbf{d}$ to each $\mathbf{d}_{s}$. This results in $N$ rays at the same distance to the surface as the original ray, with MipNeRF Gaussians [5] aligned along ray directions.
Using these virtual rays, we construct sampling batches at positions $\mathbf{x}$, and directions $\mathbf{d}$ to employ surface light field regularisation. Both batches inherit covariances from the virtual rays $\mathbf{\Sigma}_{s}$ for anti-aliasing as in ZipNeRF [7].
Figure 3: A visual depiction of the two sampling batches localised at a surface in the scene. After defining a ray surface intersection $\mathbf{x}_{\star}$, we use uniform samples on a unit ball to sample the surface through multiple viewing angles, and the local 3D volume.
Directional Sampling: Specularities at a surface are piecewise smooth, taking on the characteristics of what is being reflected. By sampling a point on a surface in a range of viewing directions and characterising the colour distribution, we can quantify how closely it matches the surface light field model introduced above. The directional batch $(\mathbf{x}_{\phi},\mathbf{d}_{\phi})$ samples at a single spatial location, but through a range of outward viewing angles:
$\quantity(\mathbf{x}_{\phi},\,\mathbf{d}_{\phi})=\quantity(\mathbf{x}_{\star},\,\mathrm{sgn}(\mathbf{d}_{s}\cdot\mathbf{n}_{\star})\mathbf{d}_{s}),$ (2)
where $\mathrm{sgn}(\cdot)$ is the sign function.
Spatial Sampling: In a NeRF, geometry is characterised by the density field $\tau(\mathbf{x})$, whose gradient with respect to position $\mathbf{x}$ approximates surface normals $\mathbf{n}\approx-\nabla_{\mathbf{x}}\tau$. By sampling density and normals of neighbouring points, and enforcing consistency between these values we can provide additional geometric supervision towards a surface model. The spatial batch $(\mathbf{x}_{m},\mathbf{d}_{m})$ consists of points $\mathbf{p}$ in the unit ball located at the surface $\mathbf{x}_{\star}$,
$(\mathbf{x}_{m},\,\mathbf{d}_{m})=\quantity(\mathbf{x}_{\star}+\sigma_{\star,r}\mathbf{p}\sqrt{2\ln{2}},\,\mathbf{d}_{s}),$ (3)
where $\sigma_{\star,r}$ is the radial variance of the sample as per mip-NeRF [4]. This samples the NeRF within a single pixel’s conical frustum, ensuring regularisation occurs at the scale of the training data; images closer to the scene regularise at finer detail. Our sample volume adapts in scale towards a minimum as the NeRF localises density during training during the later stages of training, as shown in the supplement. Since the colours go unused, the spatial batch uses the directions $\mathbf{d}_{s}$ of the virtual rays.
3.4 Local Smoothness
We regularise the geometry of the representation by penalising sparse and irregular density, allowing for a smooth and continuous sheet-like density field to form. These constraints are valid except where topological edges or corners exist; we therefore formulate $L_{1}$-norm regularisation terms to allow for these local features to form when required.
The spatial sampling batch encompasses the region in front of and behind the candidate surface. Using the normal at the candidate surface, we penalise points proportional to their density $\tau_{j}$ and perpendicular distance from the surface encouraging a plane of density to form perpendicular to the normal. Points far from the surface should have low density, and those on the same plane as the sample point should be dense. This loss term,
$\mathcal{L}_{d}=\lambda_{d}\sum_{\mathbf{r}\in\mathbf{R}}w_{\star}\sum_{j=0}^{N-1}\quantity(1-e^{-\tau_{j}})\quantity|\frac{(\mathbf{x}_{m,j}-\mathbf{{x}}_{\star})}{\norm{(\mathbf{x}_{m,j}-\mathbf{{x}}_{\star})}}\cdot\hat{\mathbf{n}}_{\star}|,$ (4)
sums over $N$ samples in the spatial batch and is also weighted by the surface point $w_{\star}$ rendering weight. This ensures that early during training, the NeRF is able to change surface locations as more rays sample the scene. We do not place a stop gradient on $\hat{\mathbf{n}}_{\star}$, allowing for the surface to reorient based on the density of samples in the local volume.
Normals in this local region should exhibit similar smoothness, and so the term,
$\mathcal{L}_{n}=\lambda_{n}\sum_{\mathbf{r}\in\mathbf{R}}w_{\star}\sum_{j=0}^{N-1}\quantity(1-e^{-\tau_{j}})\frac{\quantity(1-\quantity(\hat{\mathbf{n}}_{j}\cdot\hat{\mathbf{n}}_{\star}))}{2},$ (5)
is imposed proportional to the angle between the normal at $\mathbf{x}_{\star}$ and those at each sample point $\mathbf{p}_{j}$. This term encourages dense samples to have a normal which is parallel with that at the surface point. Both terms localise density and encourage smooth, continuous geometries. We weight these terms by hyperparameters $\lambda_{d}$ and $\lambda_{n}$ respectively.
3.5 Visually Realistic Appearance Separation
Although highly specular surfaces or highly Lambertian surfaces are well represented with a reflection encoding, surfaces with both a strong Lambertian and specular component exhibit shape-radiance ambiguity. This is characterised by incorrect geometry with a view-dependent colour bias; that is the minimum specular colour over all viewing angles is non-zero to satisfy the appearance in all training view-points. More details are found in the supplementary.
To encourage the surface light field separation of the Lambertian and view-dependent terms we propose two additional losses based on directionally sampling a point. To encourage a minimum specular bias, we penalise the normalised specular colour over viewing angles in the directional sampling batch $\mathbf{c}_{s,\phi}$,
$\mathcal{L}_{b}=\lambda_{b}\sum_{\mathbf{r}\in\mathbf{R}}w_{\star}\sum_{j=0}^{N-1}\norm{\frac{\mathbf{c}_{s,j}}{\mathrm{sg}\quantity(\norm{\max_{\phi}\,\mathbf{c}_{s}})}}^{2},$ (6)
where $\mathrm{sg}(\cdot)$ is the stop gradient operator. Where the specular distribution is uniform over viewing angles, this loss term is maximal. This ensures that the appearance of a point must accumulate most of its signal in either the specular or the Lambertian term - not in both. To maintain a piecewise smooth distribution in the specular, we enforce a total variation loss over the specular colour:
$\mathcal{L}_{s}=\lambda_{s}\sum_{\mathbf{r}\in\mathbf{R}}w_{\star}\sum_{j=0}^{N-1}\mathrm{STV}_{j}(\mathbf{d}_{\phi},\mathbf{c}_{s}).$ (7)
This total variation is applied over the surface of the sampling sphere as a form of graph total variation. This term,
$\mathrm{STV}_{j}(\mathbf{d}_{\phi},\mathbf{c}_{s})=\sum_{k}\frac{1}{2}(\mathbf{d}_{j}\cdot\mathbf{d}_{k}+1)\norm{\mathbf{c}_{j}-\mathbf{c}_{k}}_{1},$ (8)
is based upon the $k$-Nearest Neighbours of each sample point and weighted by the cosine distance between the sample and its neighbours. This approach provides a localised consistency which is edge-aware continuous over the sampling sphere, capturing occlusions. Both of these terms smooth the specular distribution and increase the reliance on the diffuse component of the model. These are weighted by hyperparameters $\lambda_{b}$ and $\lambda_{s}$ respectively.
3.6 Curriculum Learning
We impose our surface regularisation under curriculum learning to maintain visual fidelity and improve geometry. As training progresses and additional regularisation is imposed more frequently, the representation is refined to satisfy the geometric and appearance characteristics of a surface light field. The surface constraints presented in Sections 3.4-3.5 are of similar computational complexity to a regular training step, requiring a second pass through both MLPs and therefore taking close to double the time. By using curriculum learning we trade off the improvement to the representation with training time.
We schedule regularisation to occur in a staircase schedule of powers of 2: every 512 iterations early in the training, down to every 4 iterations in the final stages. On this schedule, there is approximately a 25% increase to training time.
For each regularisation sample, we add on each regularisation loss $\mathcal{L}$ to the losses for the combined ZipNeRF and Ref-NeRF model structure. More model implementation details are provided in the supplementary material, along with hyperparameter weights.
4 Results
Our regularisation is implemented in JAX [13] and based on the ZipNeRF codebase [7]. We utilise the improvements to quality and speed in ZipNeRF combined with the physics-based model structure introduced in Ref-NeRF. We compare to prior works MipNeRF360 [5] and Ref-NeRF [43].
We evaluate Surf-NeRF on the Shiny Real and the Shiny Objects [43] datasets used in Ref-NeRF as the main benchmark for our approach. We also introduce a captured dataset consisting of 4 complex reflective objects under controlled illumination, called the Koala dataset. Each scene contains approximately 40 training images and 10 test images with ground truth poses obtained using a Universal Robots UR5e robotic arm. We select objects that include non-planar specularities, fine details and specularities which present as virtual images in front of the surface. Further details are provided in the supplementary material.
4.1 Shiny Objects Dataset
Table 1 summarises results on the Shiny Objects dataset. We compare visual fidelity by peak-signal-to-noise ratio (PSNR) in decibels (dB) and structural similarity scores (SSIM), and geometric fidelity by mean angular error (MAE) of surface normals in degrees and root mean squared error (RMSE) of the disparity in inverse units. We compare across positionally encoded methods (PE) based on MipNeRF360 [5] and “grid-based” methods (Grid) based on ZipNeRF [6]. This includes baselines: MipNeRF with a diffuse channel (MipNeRF+Diff), and lattice and cubic grid based ZipNeRF with reflection parameterisation (Zip+Ref-NeRF). We train the PE implementations using V2-8 TPUs, and the Grid methods on four NVIDIA V100 32GB GPUs. Scene results are included in the supplementary materials.
Table 1: Summary on the Shiny Objects [43] dataset between visual metrics (PSNR, SSIM) and geometric metrics (MAE, RMSE). Our method yields significantly more accurate normals, trading off a marginal decrease to PSNR. Yellow, orange and red are the third, second, and first scores.
Model PSNR $\uparrow$ SSIM $\uparrow$ MAE $\downarrow$ RMSE $\downarrow$
PE MipNeRF 31.29 0.943 56.17 0.125
MipNeRF+Diff 30.84 0.936 60.00 0.122
Ref-NeRF 33.21 0.971 25.89 0.117
Surf-NeRF (Reg. Only) 33.01 0.967 22.15 0.117
Grid ZipNeRF 31.02 0.952 19.47 0.209
Zip+Ref-NeRF 33.14 0.964 14.70 0.195
Lattice+Ref-NeRF 33.21 0.964 13.76 0.202
Surf-NeRF (No Lattice) 32.13 0.954 13.35 0.202
Surf-NeRF (Ours) 32.82 0.955 10.60 0.198
Across both the positional encoding based and hash based networks we see comparable performance to Ref-NeRF in visual fidelity, demonstrating our ability to maintain visual fidelity and encourage a more geometrically accurate representation. In the positionally encoded approaches we see a 14.4% improvement to normal accuracy, with notable improvements around regions of high specularity. Surf-NeRF attains a 27.9% improvement to normals for the grid based implementation. Given the discretisation of the grid-based approaches, we see comparatively higher disparity errors across the dataset but improved normals from the interpolation of encodings.
In Figure 4 we demonstrate that the proposed regularisation is able to substantially improve surface normals and the surface consistency. This also separates the Lambertian scene content, including the racing stripes as in 5. Similarly, on the toaster scene we successfully separate the Lambertian toast and retain its reflection in the specular term.
Figure 4: Geometry comparison between normals achieved by Surf-NeRF and traditional grid implementations on the car and coffee scenes. Ref-NeRF* indicates the Zip+Ref-NeRF model. Regularising the density field improves the continuity of surfaces reducing holes and divots. The inclusion of the permutohedral lattice encoding significantly improves surface normals by better approximating smooth surface curvature.
Figure 5: Appearance separation between Surf-NeRF and baseline method Zip+Ref-NeRF. Colour renderings are comparable, but Surf-NeRF produces physically based separation of Lambertian and specular scene content, such the racing stripes and toast.
4.2 Shiny Real Dataset
The visual fidelity results on the Shiny Real dataset [43] are shown in Table 2 for Surf-NeRF and comparison baseline models. We demonstrate comparable visual fidelity to other state of the art NeRF variants demonstrating no appreciable drop in visual fidelity compared to non-regularised approaches on the positional encoding based models. The complex appearances of these datasets combined with the hash encoding sees a small drop in PSNR values for grid-based variants. Surf-NeRF shows decreased reliance on floaters to explain appearance as shown in the supplement.
Table 2: Summary of numerical results for captured Shiny Real dataset [43]. Surf-NeRF maintains visual fidelity on captured scenes, improving SSIM over the PE based methods.
Model PSNR SSIM
PE MipNeRF 23.72 0.543
MipNeRF+Diff 23.62 0.543
Ref-NeRF 23.65 0.529
Surf-NeRF (Reg. Only) 23.26 0.573
Grid ZipNeRF 23.63 0.626
Zip+RefNeRF 22.63 0.626
Lattice+RefNeRF 22.94 0.614
Surf-NeRF (No Lattice) 22.48 0.600
Surf-NeRF (Ours) 23.04 0.600
4.3 Koala Dataset
Table 3 shows the strength of the proposed regularisation on complex reflective scenes. Figure 6 demonstrates the ability for the proposed regularisation to separate Lambertian and specular colours in a more physically consistent manner, and to overcome failure cases where specularities are not densely sampled by the training data. Additional results are provided in the supplementary.
Table 3: Summary of numerical results for captured Koala dataset. Surface regularisation helps scene convergence around complex specular geometry improving PSNR.
Model PSNR SSIM
ZipNeRF 27.79 0.663
Zip+RefNeRF 23.69 0.632
Surf-NeRF (No Lattice) 27.24 0.655
Surf-NeRF (Ours) 26.86 0.654
Figure 6: The “Gold Head” and “Shiny Ball” scenes from the Koala dataset. Top rows contain RGB and median depth renderings. The Gold Head scene shows complete separation of Lambertian and specular appearance for the mannequin. The Shiny Ball dataset represents a failure case for Zip+Ref-NeRF baseline, however the proposed regularisation successfully reconstructs and appropriately places the majority of the colour into the diffuse channel.
4.4 Ablation Study
As shown in Table 4 including our full approach provides strong performance across both the visual and geometric metrics on Shiny Objects [43]. Our appearance based regularisation is coupled closely to the geometry of the scene. The permutohedral lattice encoding significantly improves surface normals coupled with our regularisation.
Table 4: Ablation study on Surf-NeRF regularisation terms on the Shiny Objects dataset. All terms contribute to overall performance. $\bigstar$ ablates deterministic sphere sampling, # ablates the permutohedral lattice and all $\mathcal{L}$ terms match those in the paper.
[…]
33.03 0.962 15.47 0.181
32.55 0.959 15.81 0.192
32.83 0.961 14.83 0.200
33.32 0.965 16.34 0.196
33.12 0.964 12.21 0.183
32.76 0.958 11.40 0.196
32.96 0.964 10.90 0.199
33.14 0.969 10.74 0.201
Rand. Samples 32.26 0.954 13.70 0.203
(No Lattice) 32.13 0.954 13.35 0.202
(Ours) 32.82 0.955 10.60 0.198
Table 5: Parameter study on regularisation frequency. There is a tradeoff to regularisation frequency between quality, time and effect, as in Figure 1. The 512/4 regime balanced multi-view photometric losses with the surface light field regularisation.
Initial Freq. Final Freq. PSNR $\uparrow$ SSIM $\uparrow$ MAE $\downarrow$ RMSE $\downarrow$
1024 8 32.63 0.955 12.46 0.204
512 4 32.82 0.953 10.60 0.198
256 2 30.59 0.946 18.99 0.220
128 1 29.77 0.926 28.37 0.218
Table 5 demonstrates a parameter study on our curriculum learning approach. Regularising frequency forces the NeRF to learn better separation of Lambertian and view-dependent channels, as in Figure 1 but reduces overall geometric quality by strongly manipulating density earlier in training. See the supplementary for more detail.
4.5 Surface Regularisation as Finetuning
We demonstrate surface regularisation as a fine-tuning step, after a NeRF has been trained on pre-trained Zip+Ref-NeRF and ZipNeRF model on the car scene from the Shiny Objects dataset. Since ZipNeRF has no Lambertian colour term, we do not include $\mathcal{L}_{b}$, and apply our sphere total variation regularisation $\mathcal{L}_{s}$ to the view-dependent colour $\mathbf{c}$, demonstrating applicability to models without reflection parameterisation. The spatial sampling terms depend only on the geometry and can be applied to other density-based neural fields without modification. We continue training at a fixed learning rate of $3.25\times 10^{-4}$ and present our results in Table 6. Normals are improved over the base model and naïve approach where the model continues training without our regularisation. Finetuning with the reflection parameterisation, as in our Zip+RefNeRF model slightly increases disparity errors as the NeRF changes density placement during the additional training.
Table 6: Surf-NeRF as a finetuning step with applicability to other NeRF variants. Additional training refines surface normals.
Base Extra PSNR SSIM Median MAE RMSE
ZipNeRF - 27.34 0.932 24.21 0.220
2.5k ZipNeRF 27.32 0.931 24.30 0.220
2.5k SurfNeRF 27.33 0.931 23.57 0.218
ZipRefNeRF - 30.25 0.957 18.44 0.246
2.5k ZipRefNeRF 30.14 0.956 18.59 0.251
2.5k SurfNeRF 30.13 0.956 18.29 0.248
4.6 Limitations
Modelling the scene as surfaces, volumetric elements like hair or subsurface scattering may have their representation degraded with the proposed regularisation. Our methodology takes twice as long on each regularisation step compared to a vanilla training step owing to two passes through the network. We enforce geometric consistency during training at the cost of reducing photometric accuracy, leading to slightly reduced PSNR scores.
5 Conclusion
We have proposed Surf-NeRF, a novel regularisation approach with four physically derived regularisation terms and a lattice hash encoding, and shown these to improve geometry in a NeRF by curriculum learning towards a surface light field model. This ensures smoothly varying surface normals and geometry with more realistic Lambertian and view-dependent appearance. This enables a NeRF to adjust geometric structure without greatly sacrificing visual fidelity, improving accuracy of up to 27.8% in normals and 6.7% in disparity. This work is a key step towards better radiance-based representations for geometry critical tasks.